OpenStudy (anonymous):

Help me solve this using the substitution method? -3x + 8 = -1 & 4x - y = 11

OpenStudy (anonymous):

please explain this to me step by step

OpenStudy (arhin):

i can help

OpenStudy (anonymous):

Can you?

OpenStudy (arhin):

yes but the ques is not clear

OpenStudy (anonymous):

it is -3x+8=-1 4x - y = 11 and i need to solve this using the substitution method

OpenStudy (arhin):

the 1st eqn is it -3+8=-1 or -3+8y=-1

OpenStudy (anonymous):

-3x+8y=-1

OpenStudy (anonymous):

sorry.

OpenStudy (arhin):

ok i can help

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

does y = 4x - 11?

OpenStudy (arhin):

make y the subject in the 1st eqn y=4x-11 then substitute it into the 2nd eqn -3x-(4x-11)=11 then x=3 substitute ur answer into any of the eqns to get y=1

OpenStudy (anonymous):

oh okay! I did it :) thank you. can you please help me on this one as well? 12x - 2y = 0 -7x + y = -1

OpenStudy (arhin):

yes but u have to be my fun and a medal

OpenStudy (anonymous):

of course.

OpenStudy (arhin):

10x alot

OpenStudy (anonymous):

np..can u help me?

OpenStudy (arhin):

yes i can

OpenStudy (anonymous):

ok:)

OpenStudy (arhin):

make y the subject in eqn 1 y=6x substitute y=6x into eqn 2 -7x+6x=-1 solve for x x=1 and then solve for y by substituting y=1 into any of the eqns u will get y=6

OpenStudy (anonymous):

okay, thank you. you're a great help.. I'm just wondering on this equation.. 3x + 12y = 24 x - 5y = 17 will x = -1?

OpenStudy (arhin):

no rather y=-1

OpenStudy (anonymous):

oh ok, thank you so much :)

OpenStudy (arhin):

u are welcome, do have any more questions

OpenStudy (anonymous):

no, that's it. thank you!

OpenStudy (arhin):

then lets have a chat

OpenStudy (anonymous):

sure, about what?

OpenStudy (arhin):

about each other country,age,college etc

OpenStudy (anonymous):

are you a girl or guy?

OpenStudy (anonymous):

& you can message me if u want..